If l, m be the portions of the axes of x and y intercepted by the tangent at any point ( x,y) on the curve ( x/a )^2/3+  (  y/b)^2/3  , show that ( l^2/a^2) +(m^2/b^2)=1


1
Expert's answer
2021-12-01T13:48:36-0500

l, m be the portions of the axes of x and y intercepted by the tangent at any point ( x,y) on the curve

"( \\frac{x}{a} )\ufeff^{\\frac{2}{3}}+ (\\frac{y}{b})\ufeff^{\\frac{2}{3}}=1"

"b^{\frac{2}{3}}x^{\frac{2}{3}}+a^{\frac{2}{3}}y^{\frac{2}{3}}=a^{\frac{2}{3}}b^{\frac{2}{3}}" ••••equation(1)

The parametric coordinate of the point (x,y) is ("acos³\theta , bsin³\theta)" , "\theta" is a parameter.

By the problem, equation of tangent is

"\frac{x}{l}+\frac{y}{m}=1" ••••equation(2) to the curve at ( acos³"\theta" , bsin³"\theta)"

Now we find equation of tangent at (acos³"\theta" , bsin³"\theta)"

Differentiating equation (1) with respect to x we get

"\frac{2}{3}b^{\frac{2}{3}}x^{-\frac{1}{3}} + \frac{2}{3}a^{\frac{2}{3}}y^{-\frac{1}{3}} \frac{dy}{dx}=0"

=> "\frac{dy}{dx} = -\frac{ {b^{\frac{2}{3}}y^{\frac{1}{3}}}}{ {a^{\frac{2}{3}}x^{\frac{1}{3}}}}"

Now "[\frac{dy}{dx}]_{(acos³\theta , bsin³\theta)}= -\frac{ {b^{\frac{2}{3}}(bsin³\theta)^{\frac{1}{3}}}}{ {a^{\frac{2}{3}}(acos³\theta)^ {\frac{1}{3}}}} =-\frac{b}{a}tan\theta"

So equation of tangent at "(acos³\theta , bsin³\theta)" will be

"y-bsin³\theta" ="-\frac{b}{a}tan\theta(x-acos³\theta)"

=> "(acos\theta)y + (bsin\theta)x = absin\theta cos³\theta + abcos\theta sin³\theta"

=> "(bsin\theta)x +(acos\theta)y = absin\theta cos\theta (cos²\theta + sin²\theta)"

=> "(bsin\theta)x +(acos\theta)y = absin\theta cos\theta"

=> "\frac{(bsin\theta)x +(acos\theta)y}{absin\theta cos\theta }= 1"

=> "\frac{x}{acos\theta}+\frac{y}{bsin\theta}=1" ••••equation(3)

Obviously equation (2) and equation (3) are identical.

So equating coefficients we get

"\frac{l}{acos\theta}=\frac{m}{bsin\theta}=\frac{1}{1}"

=> l = acos"\theta" and m = bsin"\theta"

=> "\frac{l}{a} = cos\theta" and "\frac{m}{b} = sin\theta"

So "{(\frac{l}{a})}^{2} +{(\frac{m}{b})}^{2}= cos²\theta + sin²\theta"

=> "{(\frac{l}{a})}^{2} +{(\frac{m}{b})}^{2}= 1"


=> "{(\frac{l²}{a²})}+{(\frac{m²}{b²})}=1"

Hence proved.